Please help me solve this age problem:
Lanny and Runiel where born on the same day.
Lanny is 16 years older than Runiel.
In 1998, he is 5 times older.
In what year will she be twice as old?

Answer by ankor@dixie-net.com(12695)

:
L is 16 years older than R.
L = R + 16
:
In 1998, he is 5 times older.
2008 - 1998 = 10 yrs ago
:
(L-10) = 5(R-10)
L - 10 = 5R - 50
L = 5R - 50 + 10
L = 5R - 40
Find out how old they are now:
Substitute (R+16) for L, solve for R
R + 16 = 5R - 40
16 + 40 = 5R - R
56 = 4R
R = 56/4
R = 14 yrs old now
then
14 + 16 = 30 yrs old is L now
:
In what year will she twice as old?
Let x = no. of yrs from now she will be twice as old:
L + x = 2(R+x)
Substitute their ages in the above equation
30 + x = 2(14+x)
30 + x = 28 + 2x
30 - 28 = 2x - x
x = 2 yrs; 2010 L is twice as old as R
:
This is true in 2010 (2 yrs from now) L will be 32 and R will be 16,